OpenStudy (anonymous):

There are too many steps can someone write it out slowly? [a-4{2a-3(a-2)}]

OpenStudy (anonymous):

Work from the inner parentheses outwards: multiply: 3(a-2) first what do u get?

OpenStudy (anonymous):

3a-6

OpenStudy (anonymous):

good then simplify: {2a-(3a-6)}. What do u get?

OpenStudy (anonymous):

6a and 12a ?

OpenStudy (anonymous):

2a-3a+6 = ?

OpenStudy (anonymous):

a-4(2a-3a+6)=a-8a+12a-24=5a-24

OpenStudy (anonymous):

-a+6 is that it

OpenStudy (anonymous):

you then: a-4{-a+6}, distribute the -4 to the two terms in the parentheses by a-4*(-a) -4*(6), then simplify from here.

OpenStudy (anonymous):

thank you all

OpenStudy (anonymous):

Sure. Follow these steps n u can't go wrong always work from inside out.